OpenStudy (unknownrandom):

The ratio of 3 more than a number to the square of 1 more than that number is less than 1. Find the numbers which satisfy this statement.

OpenStudy (unknownrandom):

How do I set this up?

OpenStudy (anonymous):

let the number be \(a\) then \[\frac{3+a}{(a+1)^2}<1\]

OpenStudy (unknownrandom):

I see now. I have been on this homework for too long. Thanks!

OpenStudy (anonymous):

\[3+a<a^2+2a+1\] \[a^2+a-2=(a+2)(a-1)>0\] s0 \[a>1,a<-2\]

OpenStudy (mathmale):

There are two "critical numbers" here. One is the value of ' a ' for which the denominator is zero. The other is the value of ' a ' for which the numerator is zero. What are these two "critical numbers?)?

OpenStudy (unknownrandom):

It looks like -1 and -3

OpenStudy (mathmale):

I agree. Draw a number line and mark -1 and -3 on that line. These two critical values will divide your number line into 3 intervals. Would you please identify these intervals. 

OpenStudy (mathmale):

I trust you know you can copy my graph and add to it.

OpenStudy (unknownrandom):